# Poincare field theory for massless particles

### v1

@@ 9 Conclusion
Our systematic approach has verified Weinberg observation [4] that Lorentz invariance is violated with terms in the form of local gauge transformations. The source of this effect is non commutativity of components of Pauli-Lubanski vector. We check this statements for vector fields and for second and fourth rank tensor fields.
Let us stress that helicity is crucial for determining field equations for irreducible components of some field $\Psi^{A}$ and to find the number of theirs degrees of freedom. Note that we can describe the same theory using different fields. For example, as it shows Eq.(8) all representations with helicities $\lambda=\pm 1$ (vector field, symmetric and antisymmetric second rang tensor) have one more vector $\breve{p}^{a}_{\pm}$ than $\breve{p}^{a}_{\mp}$ and describe electromagnetic interaction. Also, all representations with helicities $\lambda=\pm 2$ (symmetric second rang tensor and fourth rang tensor with appropriate symmetries) have two more vector $\breve{p}^{a}_{\pm}$ than $\breve{p}^{a}_{\mp}$ and describe gravitation interaction. Consequently, contrary to the common belief that only highes helicity of certain tensor is relevant, we claim that all helicities are relevant. Different fields with the same helicity can describe the same theory, by sometimes it not easy to write field equation in terms of that fields.
It should be emphasize the important logical sequence in the present article Covariant definition of helicity $\to$ little group invariance $\to$ local gauge transformations $\to$ equations of motion for massless fields. In other words, the necessity of existence of local gauge transformations for massless fields originates from requirement for Lorentz invariance.

### v2

@@ 12 Conclusion
Our systematic approach has verified Weinberg observation [4] that Lorentz invariance is violated with terms in the form of local gauge transformations. The source of this effect is non commutativity of components of Pauli-Lubanski vector. We check this statements for tensor fields with arbitrary rank.
Let us stress that helicity is crucial for determining field equations for irreducible components of some field $\Psi^{A}$ and to find the number of theirs degrees of freedom. Note that we can describe the same theory using different fields. For example, as it shows Eq.(8) all representations with helicities $\lambda=\pm 1$ (vector field, symmetric and antisymmetric second rank tensor) have one more vector $\breve{p}^{a}_{\pm}$ than $\breve{p}^{a}_{\mp}$ and describe electromagnetic interaction. Also, all representations with helicities $\lambda=\pm 2$ (symmetric second rank tensor and fourth rank tensor with appropriate symmetries) have two more vector $\breve{p}^{a}_{\pm}$ than $\breve{p}^{a}_{\mp}$ and describe gravitation interaction. Consequently, contrary to the common belief that only highes helicity of certain tensor is relevant, we claim that all helicities are relevant. Different fields with the same helicity can describe the same theory, but sometimes it not easy to write field equation in terms of that fields.
Although the main contribution of the present article is proof that principle field equations (8) produce all known free field theory equations, this paper is also improvement over the set of articles with highest helicity [5, 6, 7, 8], well known in literature. For details and relations with these papers see Section $11$.
It should be emphasize the important logical sequence in the present article Covariant definition of helicity $\to$ little group invariance $\to$ local gauge transformations $\to$ equations of motion $\to$ actions for massless fields. In other words, the necessity of existence of local gauge transformations for massless fields originates from requirement for Lorentz invariance.